# GAPS: Guiding Dynamic Android Analysis with Static Path Synthesis

### v1

@@ 8. Conclusions
In this paper, we present GAPS, a tool aimed at path reconstruction through the generation of Android apps’ call graph, that also offers a module to traverse the identified paths automatically. In our evaluation, we show that GAPS can reconstruct at least one path for 88.24% target methods and traverse 57.44% of such paths for apps belonging to the AndroTest benchmark with an average analysis time of 4.27 seconds. In contrast, state-of-the-art tools achieve significantly lower reachability: GoalExplorer (9.69%), APE (12.82%), Guardian (17.12%), FlowDroid (58.81%), DroidReach (9.48%), with longer or less targeted analysis phases. On the other hand, the results on the real-world apps dataset report that GAPS can reconstruct paths for 62.03% of the target methods, averaging an analysis time of 278.9 seconds per app, and dynamically proving the reachability for 59.86% of them.
Overall, we can conclude that GAPS can fill the gap between static and dynamic analyses, guiding a targeted app exploration and solving the method reachability problem.

### v2

@@ 8. Conclusions
In this paper, we present GAPS, a framework designed for statically reconstructing paths that guide the dynamic analysis until the target method is executed. In our evaluation, we show that GAPS can reconstruct at least one path for 88.24% of the target methods and traverse 57.44% of such paths in apps belonging to the AndroTest benchmark, with an average analysis time of 4.27 seconds. In contrast, state-of-the-art tools achieve significantly lower performance in static path reconstruction and dynamic method reachability: GoalExplorer (9.69%), APE (12.82%), Guardian (17.12%), FlowDroid (58.81%), DroidReach (9.48%), with longer or less targeted analysis phases. On the other hand, the results on the real-world apps dataset report that GAPS can reconstruct paths for 62.03% of the target methods, averaging an analysis time of 278.9 seconds per app, and dynamically executing 59.86% of them.
Overall, we can conclude that GAPS can bridge the gap between static and dynamic analyses, guiding targeted app exploration and addressing the method reachability problem.